# Keyword-driven Retrieval-Augmented Large Language Models for Cold-start User Recommendations

### v1

@@ 5. Discussion and Conclusion
In this work, we investigate the capabilities of Augmented LLMs for cold-start user recommendations by leveraging keywords extracted from user reviews. Our experimental results demonstrate that the candidates retrieval and LLM-based candidates ranking methods proposed in KALM4REC effectively address cold-start issues. For retrieving potential restaurant candidates, we present CBR and MPG, keyword-based methods, which outperform conventional approaches. We then employ LLMs to re-rank the obtained candidates using designed prompting strategies that incorporate keywords to represent users and restaurants. Comprehensive experiments indicate that KALM4REC is capable of handling cold-start user scenarios. Additionally, our framework shows the potential of integrating different retrieval and language models to achieve promising performance in the future.
We acknowledge that certain factors may affect the performance of our work. For instance, while the constructed graph in MPG is effective, the connections between keyword nodes should also be considered, as similar keywords may need to pass information to each other before reaching the restaurant nodes. However, implementing this could be computationally costly due to the vast number of keywords in large cities like London. Additionally, the number of keywords introduced into the prompt may be reduced by increasing the number of example introductions, as shown in the experiments. For future work, we plan to develop technical approaches to enhance the robustness of each stage in our framework, leveraging the insights drawn from our comprehensive set of experiments.

### v2

@@ 5. Discussion and Conclusion
In this work, we investigate the idea of augmenting LLMs for cold-start user recommendations with keywords extracted from user reviews. For retrieving potential candidates, we present CBR and MPG, keyword-based methods, which outperform conventional approaches. We then employ LLMs to re-rank the obtained candidates using designed prompting strategies that incorporate keywords to represent users and items. Comprehensive experiments indicate that KALM4REC is capable of handling cold-start user scenarios effectively. Additionally, our framework shows the potential of integrating different retrieval and language models to achieve promising performance in the future for multiple domains. Several factors may affect the performance of our framework. For instance, while the constructed graph in MPG is effective, the connections between keyword nodes should also be considered, as similar keywords may need to pass information to each other before reaching the restaurant nodes. However, implementing this could be computationally costly when the number of keywords is large. Additionally, the number of keywords introduced into the prompt may be reduced by increasing the number of example introductions, as shown in the experiments. For future work, we plan to develop technical approaches to enhance the robustness of each stage in our framework, leveraging the insights drawn from our comprehensive set of experiments.